Please Help:
Find the interest. Round to the nearest cent.
$1980 at 15% for 5.5 years
A) Interest = $726.00
B) Interest = $54.00
C) Interest = $163.35
D) Interest = $1633.50

Answer by rfer(16322)

1980*.15=297*5.5=1633.50 (D)
find the interest per year times the number of years